My Thesis Title

My Full Name

A subthesis submitted in partial fulfillment of the degree of

Bachelor of Science (Honours) at The Department of Computer Science

Australian National University

September 1997

\begin{matrix}c \\ ◯ \end{matrix} My Full Name

Typeset in Palatino by TEX and L^{A}TEX 2ε.

Except where otherwise indicated, this thesis is my own original work. My Full Name 5 September 1997 To fluffy the cat.

Acknowledgements

Here is where you thank the people who helped you along the way…

Abstract

This should be the abstract to your thesis… x

Contents

Acknowledgements vii Abstract ix 1 An Introduction to My Thesis 1

1.1 The Basis for this Work . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1.1 A Theoretical Framework . . . . . . . . . . . . . . . . . . . . . . 1 2 The First Experiment 3

2.1 Why I Did It . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 What I Did . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3 The Second Experiment 7 3.1 Why I Did It . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.2 What I Did . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

4 Conclusion 9

4.1 Why this is a Very Clever Thesis . . . . . . . . . . . . . . . . . . . . . . 9 A Some Other Stuff 11

A.1 Why I Did It . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 B More Stuff 13 Bibliography 15 Index 16 xii Contents

Chapter 1

An Introduction to My Thesis

1.1 The Basis for this Work

1.1.1 A Theoretical Framework

2 An Introduction to My Thesis

Chapter 2

The First Experiment

2.1 Why I Did It

2.2 What I Did

We add a few blank pages here to illustrate section headings on odd pages (other than the first page in a new chapter).

Also, here is an example of a citation [Lamport 1994], and another one [Knuth 1986], and another [Goossens et al. 1994]. If in the context, it makes sense to talk about the work of an author in a more integrated way, like Lamport [1994], that can be done too. Because the information on this page is so important, it has been indexed too (see the index at the back of this thesis).^{1} 4 The First Experiment

1However, it is strongly advisable to leave indexing until the thesis is complete. Donald Knuth says to allow about a day for the task of indexing. In my experience he was spot-on.

Note the chapter name at the top of this even (lefthand) page…

§2.2 What I Did 5

Note the section heading at the top of this odd (righthand) page… 6 The First Experiment

Chapter 3

The Second Experiment

3.1 Why I Did It 3.2 What I Did

8 The Second Experiment

Chapter 4

Conclusion

4.1 Why this is a Very Clever Thesis

10 Conclusion

Appendix A

Some Other Stuff

A.1 Why I Did It

12 Some Other Stuff

Appendix B More Stuff

13 14 More Stuff

Bibliography

Goossens, M., Mittelbach, F., and Samarin, A. 1994. The L^{A}TEXCompanion. Addison Wesley. (p. 3)

Knuth, D. E. 1986. The TEXbook, Volume A of Computers and Typesetting. Addison Wesley. (p. 3)

Lamport, L. 1994. L^{A}TEX: A Document Preparation System—User’s Guide and Reference Manual (2nd ed.). Addison Wesley. (p. 3)

Index

Books to read, 3 Section headings, 3